# Causal Concept Graph Models: Beyond Causal Opacity in Deep Learning

### v4

@@ 1 Introduction
Although CBMs have been shown to be semantically transparent (Kim et al., 2018; Facchini & Termine, 2021) [^1] Here by semantic transparent we mean a model that focuses on features that are meaningful and understandable by human experts in the model’s application domain., in this work we argue that these models still rely on two unrealistic assumptions (visualised in Fig. 1(b)). First, they assume that concepts are causally independent, an assumption that breaks in most real-world tasks. For instance, CBMs might treat “smoking” and other health conditions as independent predictors of an insurance premium, ignoring how intervening on smoking habits might have a significant impact on health conditions. Second, and perhaps more importantly, CBMs operate with the underlying assumption that concepts are the direct causes of the model’s predictions. These assumptions oversimplify and constrain the model’s inference structure, preventing the model from using more complex and effective concept dependencies (Beckers, 2022). As a result, the problem of causal opacity still represents a key open challenge at the intersection of DL, causality, and XAI.
Contributions
In response to this challenge, in this paper we introduce Causal Concept Graph Models (Causal CGMs, see Figure 1(c)). Causal CGMs are a new a concept-based architecture whose decision-making process is causally transparent by design, avoiding the unrealistic assumption that concepts must be causally independent and direct causes of the class prediction. The results of our experiments show that Causal CGMs can: (i) match the generalisation performance of state-of-the-art causally-opaque models, (ii) enable human-in-the-loop corrections of mispredicted intermediate reasoning steps, boosting not just downstream accuracy after corrections but also the accuracy of the explanation provided for a specific instance, and (iii) support the analysis of both interventional and counterfactual questions, thereby improving the model’s causal interpretability and supporting the effective verification of its reliability and fairness.
@@ 2 Background
A causal model (CM) is a mathematical representation of the mechanisms (rules, laws) that explain how different variables of a given target-system causally influence each other. In computer science and statistics, the standard framework for causal modelling is that of Structural Causal Models (SCM) proposed by Pearl (2009). Formally, an SCM $\mathcal{M}$ is a triplet $(\mathcal{U},\mathcal{V},\mathcal{F}_{\theta})$ where: $\mathcal{U}$ is a set of exogenous variables representing latent factors with a causal influence on the modelled target-system; $\mathcal{V}$ is a set of endogenous variables representing observable and measurable variables; $\mathcal{F}_{\theta}$ is a set of functions (or structural equations) parametrized by $\theta$ describing the causal mechanisms, which determine the values of each endogenous variable $v_{i}\in\mathcal{V}$ by computing the conditional probability $p(v_{i}|u_{i},\text{pa}(v_{i});\theta)$, where $u_{i}\in\mathcal{U}$ and $\text{pa}(v_{i})$ are, respectively, the set of exogenous and endogenous variables causally affecting $v_{i}\in\mathcal{V}$. Every SCM can be associated with a directed acyclic graph (DAG) whose nodes represent variables and edges represent direct causal connections. Interventions can be modelled in SCMs through the do-operator (Pearl, 1995; 2009) . This operator induces a modification on the model’s structure by changing the value of an endogenous variable which is fixed to a value $\kappa\in\mathbb{R}$ by external means, effectively breaking its original causal dependencies within the model. Formally, applying $do(v_{i}=\kappa)$ to a model $\mathcal{M}$ results in a new model $\mathcal{M}_{v_{i}=\kappa}$ which is identical to $\mathcal{M}$ except that the equation for $v_{i}$ is replaced with a constant value $\kappa$, removing all arrows into $v_{i}$ in the model annexed DAG.
Concept-based models
@@
Concept-based models (Kim et al., 2018; Chen et al., 2020; Yeh et al., 2020; Kim et al., 2023; Barbiero et al., 2023; Oikarinen et al., 2023) are interpretable architectures that explain their predictions using high-level units of information (i.e., “concepts”). Most of these approaches can be formulated as a Concept Bottleneck Model (CBM) (Koh et al., 2020). Given a sample’s raw features $\mathbf{x}\in X\subseteq\mathbb{R}^{d}$ (e.g., an image’s pixels), a set of $r$ concepts $c_{i}\in C\subseteq\{0,1\}^{r}$ (e.g., “red”, “round”), and a set of $l$ class labels $y_{j}\in Y\subseteq\{0,1\}^{l}$ (e.g., labels “apple” or “tomato”), a CBM estimates the conditional distribution $\prod_{j}p(y_{j}\mid\text{pa}(y_{j})=c_{1},\dots,c_{r};\theta_{f})\prod_{i}p(c_{i}\mid\mathbf{x};\theta_{g})$ where $p(c_{i}\mid\mathbf{x};\theta_{g})$ is a set of independent Bernoulli distributions. This formulation emphasises two key CBMs’ assumptions: i) concepts are direct causes of class predictions, and ii) concepts are conditionally independent of one another. Usually, concept-based models represent a concept $c_{i}$ using its predicted probability $\hat{c}_{i}\in[0,1]$. This representation, however, might significantly degrade task accuracy when the provided concepts are incomplete (Mahinpei et al., 2021). To overcome this issue, Concept Embedding Models (CEMs) (Espinosa Zarlenga et al., 2022) use high-dimensional embeddings $\mathbf{\hat{c}}_{i}\in\mathbb{R}^{z}$ to represent concepts alongside their truth degrees $\hat{c}_{i}\in[0,1]$.
@@ 3 Causal Concept Graph Models
Problem: This work addresses causal opacity denotes the difficulty in understanding the hidden causal structure behind a DL model’s inference and decision-making process. Research question: To overcome this, we aim to design DL models where each prediction can be traced back to a chain of semantically meaningful causes. Challenged assumptions: In pursuing this goal, we challenge two key assumptions of Concept Bottleneck Models (CBMs): i) humans know a priori which interpretable variables (i.e., concepts) are direct causes of a given class label (and the dataset embeds this prior knowledge), and ii) all interpretable variables used for explanations (i.e., concepts) are conditionally independent. Solution: We introduce Causal Concept Graph Models (Causal CGMs), a class of models making class predictions through cause-effect chains of interpretable variables. In what follows, we first introduce Causal CGMs and then proceed to show how their decision-making process is causally transparent. Finally, we conclude this section by describing their learning process.
@@ 3.1 Blueprint
Dropping the assumptions of CBMs—that concepts are independent and direct causes of class labels—requires a model able to capture all possible dependencies [^2] This naive formulation would lead to a cyclic probabilistic model, which becomes computationally intractable due to the difficulty in handling cycles in probabilistic models. Causal Concept Graph Models tackle this issue by ensuring the learning process is efficient and produces simple explanations by design (as described in Sec. 3.3). between interpretable variables (i.e., any classification label). To model such dependencies, Concept Graph Models duplicate the layer of interpretable variables $\mathcal{V}$ by introducing an additional layer of identical copies $\mathcal{V}^{\prime}$ [^3] A detailed analysis of this transformation, including the “unfolding” of the original cyclic model, is provided in App. B and a neural parametrization in App. D.. Using this additional layer, Causal CGMs make predictions of each $v_{i}\in\mathcal{V}$ using as possible inputs all $v_{j}^{\prime}\in\mathcal{V}^{\prime}$ for all $j\neq i$. This inference structure enables CGMs to trace each prediction in $\mathcal{V}$ back to a set of interpretable variables in $\mathcal{V}^{\prime}$. As a result, the set of copies $\mathcal{V}^{\prime}$ in a Causal CGM play the role of “explaining variables”, akin to a CBM’s concepts, while the set of variables $\mathcal{V}$ play the role of “explained variables”, akin to a CBM’s tasks.
@@ 3.2 Causal transparency in Concept Graph Models
The above theorem shows that CGM’s bipartite graph $\mathcal{G}^{\prime}$ between nodes $\mathcal{V}^{\prime}$ and $\mathcal{V}$ can efficiently represent arbitrary long cause-effect chains among interpretable variables (example in Fig. 2). While during training CGMs exploit the efficient bipartite representation, at test time CGMs can be unfolded following the graph as described in Theorem 3.2 (Sec. 3.3 describes how to learn sparse and acyclic dependencies to make the test time unfolding more efficient and interpretable; we provide more details in App. H.6). This property makes CGMs causally transparent by design, as they can make class predictions through cause-effect chains of endogenous variables, allowing humans to trace each prediction back to its underlying causes within an interpretable parental subgraph. Notice how—unlike CBMs—Causal CGMs do not require to know in advance cause (concept)-effect (task) relationships, nor require that all causes (concepts) are independent from each other, as cause-effect dependencies among interpretable variables are not given, but learnt during training.
@@ 3.3 Training Causal CGMs
Learning sparse, directed, and acyclic structures
In order to construct simple explanations, we can make Causal CGMs’ graphs sparse and acyclic. To this end, we parametrize the bipartite graph $\mathcal{G}^{\prime}$ with an adjacency matrix of learnable weights $M\in\mathbb{R}^{k\times k}$. We initialise the weights $m_{ij}\in M$ based on the conditional entropy between labels $v_{i}$ and $v_{j}$ in order to account for asymmetric concept dependencies (for other initialisation strategies, see App. F). These weights are then fine-tuned through an end-to-end learning process within the Causal CGM framework following common causal priors, where causal graphs are assumed to be sparse, directed, and acyclic (forming a Directed Acyclic Graph or DAG). We introduce a parameter $\gamma\in\mathbb{R}$ to eliminate less significant dependencies and a loss function, as described by Yang et al. (2020), to enforce the sparsity and acyclicity of the causal graph, ensuring that the adjacency matrix $A$ effectively represents a DAG:
@@
Optimisation problem
We can now state the general learning objective for Causal CGMs. Given (1) a set of entities represented by their feature vectors $\mathbf{x}\in\mathcal{X}\subseteq X$ (i.e. the input) and (2) a set of annotations for each exogenous variable $v\in\mathcal{V}\subseteq V$ (i.e. the labels), we wish to find functions $\zeta$, $s$, $f$, together with the adjacency matrix $A$, that maximise the log-likelihood of $v,v^{\prime}$, while observing $x$ (or equivalently $u=\zeta(x)$):
@@
where $\lambda_{1,2}$ are hyperparameters balancing optimisation objectives (related ablation studies can be found in App. H.3). Notice that we use the $do$-operator replacing $\hat{v}_{j}^{\prime}$ with labels $v_{j}$ to minimise leakage and provide better gradients to the endogenous predictor $f$. This enables Causal CGMs to be aware of $do$-operations during training, thus making the model effective in responding to $do$-interventions once deployed.
@@ 3.4 Causal reasoning and verification with Causal CGM
Ground-truth interventions
Causal CGMs support “ground-truth interventions” (see Figure 3(b)). Ground-truth interventions are one of the core motivations behind CBMs (Koh et al., 2020). Through ground-truth interventions, concept bottleneck models allow experts to improve a CBM’s task performance by rectifying mispredicted concepts at test time, thus significantly improving task performance within a human-in-the-loop setting. In Causal CGMs, however, ground-truth interventions have a potential impact on all endogenous variables descendant of an intervened node, which may include not only nodes corresponding to downstream tasks but also nodes corresponding to a CBM’s intermediate concepts. This enables a single concept ground-truth intervention to potentially improve the prediction of intermediate concepts as well as downstream tasks.
Causal reasoning and verification: do-interventions, counterfactuals, and blocking
Causal CGMs can answer interventional and counterfactual queries related to the model’s decision-making process using the do-operator on the unfolded SCM.
@@
Figure 3: (a) A 5-variable causal graph. (b) A ground-truth intervention fixes the error of the prediction $\hat{v_{3}}$ to the ground-truth label $v_{3}$. (c) A do-intervention sets the value of the second variable to a constant i.e., $v_{2}=0$. The intervention impacts $v_{2}$’s effects i.e., $v_{3,5}$, but does not alter $v_{2}$’s causes i.e., $v_{1}$. This operation can override ground-truth interventions. (c) A do-intervention on $v_{3}$ blocks the causal effects of $v_{2}$ on $v_{5}$. As a result, intervening on $v_{2}$ cannot alter $v_{5}$ anymore.
@@ 4.1 Key findings
Causal CGMs match the performance of causally opaque models. (Table 1)
Causal CGMs demonstrate robust generalisation across all datasets, yielding a predictive performance close to that of black-box architectures with an equivalent capacity. Causal CGMs using a pre-trained causal graph (Causal CGM+CD) tend to have slightly better label accuracy with respect to Causal CGMs where the causal graph is learned end-to-end (Causal CGM). Causal CGMs’s low variance suggests a consistent robustness with respect to weight initialisations over multiple training runs. Thanks to concept embeddings, concept incompleteness settings do not have a strong impact on Causal CGMs’ performance (results in App. H.4).
@@
Ground-truth interventions on Causal CGMs improve both concept and task accuracy as opposed to CBMs (Figures 5, 5)
In Causal CGM, the causal graph induces a natural strategy for ground-truth interventions. Indeed, the causal graph narrows down the set of variables to intervene upon: for any given node, we can just fix mispredicted labels of the node’s ancestors as intervening on other nodes will not have any impact. This property significantly decreases the required number of interventions to achieve a desired outcome (e.g., to increase a downstream task accuracy), as shown in App. H. Another advantage of Causal CGM consists in the hierarchical nature of inference which allows ground-truth interventions to impact all endogenous variables descendant of an intervened node. In particular, ground-truth interventions may affect not only nodes corresponding to downstream tasks (as in CBM and CEM), but also nodes corresponding to a CBM’s intermediate concepts. We experimentally verify this property and its impact by calculating—for nodes that were not intervened upon (including both concepts and tasks)—the change in accuracy before and after ground-truth interventions were applied on their ancestors. Our results (Figure 5 and App. H.1) show that Causal CGM improves nodes accuracy by $\sim 15$ percentage points after only 7 ground-truth interventions on CelebA. CBM and CEM, instead, achieve a similar performance only after intervening on all concepts as their architecture assumes all concepts to be mutually independent. Focusing solely on concept-level analysis, Causal CGM enhances concept accuracy by up to $\sim 25$ percentage points, compared to CBM and CEM, where a concept intervention does not impact the accuracy of other concepts (Figure 5). Causal CGM’s advantage increases with the number of concepts and connections, as a single intervention can impact a higher number of nodes in the causal graph.
@@
Causal CGMs’ endogenous predictors are causally interpretable (Figures 6(a), 6(b))
In Causal CGM the decision-making process is causally interpretable and can be analysed by visualising the learnt causal graph and structural equations as in a structural causal model, as shown in Figure 6(a) for CelebA and in Figure  6(b) for Cifar10 (more results in App. H.2). The first image shows how Causal CGM exploited known biases in CelebA to infer facial attributes. For instance, Causal CGM predicts the attribute “wearing lipstick” directly from the attribute “attractive” and indirectly from attributes such as “smiling” or “high cheek”, all attributes that are known to be strongly correlated with each other in CelebA (Ramaswamy et al., 2021; Wang & Russakovsky, 2023). Similarly, in the CIFAR-10 dataset, where concepts are automatically extracted, the model learns meaningful relationships between concepts (Figure 6(b)), such as the presence of a “port” implying the presence of a “dock”, or the connection between a “beak” and a “bird”. We can also quantify the strength of the causal dependency between two nodes by computing the probability of necessity and sufficiency (PNS) (Pearl, 2022). In the figure, we represent the PNS w.r.t. the leaf node by colouring each node with a different shade of orange (more PNS of the other datasets in App. I). This shows how, for Causal CGM, the attribute “heavy makeup” has the strongest impact on the leaf node. This high degree of causal transparency allows users to interpret Causal CGM’s inference and can be eventually exploited to identify potential biases, thereby supporting the assessment of the model’s counterfactual fairness (an example of generated counterfactuals in Table  2). As a result, users can intervene directly on the causal structure of the decision-making process and remove biases using do-interventions, as shown in the next paragraph.
@@
Causal CGMs can make two causally-related variables causally independent by blocking all paths between these variables (Table 3)
The causal transparency of the proposed approach allows users to modify the model’s decision-making process (e.g., to de-bias the model’s inference) by using do-interventions. In particular, it enables making two causally-related variables $i$ and $j$ independent by blocking all paths between the cause $i$ and the effect $j$. We experimentally verify this property and its impact by computing the Residual Concept Causal Effect i.e., the ratio between the Concept Causal Effect (CaCE) (Goyal et al., 2020) after and before blocking. The optimal value of this metric is zero, corresponding to perfect causal independence between $i$ and $j$ (i.e., an ideal de-biasing operation). The results show that, in Causal CGMs, blocking a variable in the causal graph always yields a perfect Residual Concept Causal Effect of zero across all datasets. In contrast, applying the same procedure in CBMs only reduces the average causal effect slightly, by just $3$ percentage points. CEMs not only fail to reduce the causal effect to zero but, in some cases, even increase the causal influence. These results underscore how Causal CGMs transparency enable users to manipulate the model’s decision-making process to achieve desired outcomes, as opposed to existing CBMs.
Table 3: Residual Concept Causal Effect ($\downarrow$) between causally-related variables having blocked all paths between the two variables with do-interventions on the causal graph. The optimal value is zero corresponding to perfect causal independence. Values above $100\%$ mean that the causal effect increased instead of decreasing.
checkmark dsprites celeba celeba

### v6

@@ 1 Introduction
Although CBMs are semantically transparent (Kim et al., 2018; Facchini & Termine, 2021) [^1] Here by semantic transparent we mean a model that focuses on features that are meaningful and understandable by human experts in the model’s application domain., we argue they still rely on two unrealistic assumptions (visualised in Fig. 1(b)). First, they assume that concepts are causally independent, an assumption that breaks in most real-world tasks. For instance, CBMs might treat “smoking” and other health conditions as independent predictors of an insurance premium, ignoring how intervening on smoking affects health. Second, and more importantly, CBMs assume concepts are the direct causes of model predictions. These assumptions oversimplify and constrain the model’s inference structure, preventing the model from using more complex and effective concept dependencies (Beckers, 2022). As a result, the problem of causal opacity still represents a key open challenge at the intersection of DL, causality, and XAI.
Contributions In response to this challenge, we introduce Causal Concept Graph Models (Causal CGMs, see Figure 1(c)). Causal CGMs are a new concept-based architecture whose decision-making process is causally transparent by design, avoiding the unrealistic assumption that concepts must be causally independent and direct causes of class prediction. The results of our experiments show that Causal CGMs can: (i) match the generalisation performance of state-of-the-art causally-opaque models, (ii) enable human-in-the-loop corrections of mispredicted intermediate reasoning steps, boosting both downstream accuracy and explanation correctness, and (iii) support the analysis of both interventional and counterfactual questions, improving causal interpretability and enabling effective verification of reliability and fairness. The code related to this paper is publicly available 222 https://github.com/gabriele-dominici/CausalCGM.
@@ 2 Background
Causal Models A causal model (CM) is a mathematical representation of the rules or laws that explain how different variables of a given target-system causally influence each other. In computer science and statistics, Structural Causal Models (SCM) proposed by Pearl (2009) are a key framework for causal modelling. Formally, an SCM $\mathcal{M}$ is a triplet $(\mathcal{U},\mathcal{V},\mathcal{F}_{\theta})$ where: $\mathcal{U}$ is a set of exogenous variables representing latent factors with a causal influence on the modelled target-system; $\mathcal{V}$ is a set of endogenous variables representing observable and measurable variables; $\mathcal{F}_{\theta}$ is a set of functions (or structural equations) parametrised by $\theta$ describing the causal mechanisms, which determine the values of each endogenous variable $v_{i}\in\mathcal{V}$ by computing the conditional probability $p(v_{i}|u_{i},\text{pa}(v_{i});\theta)$, where $u_{i}\in\mathcal{U}$ and $\text{pa}(v_{i})$ are, respectively, the set of exogenous and endogenous variables causally affecting $v_{i}\in\mathcal{V}$. Every SCM corresponds to a directed acyclic graph (DAG) whose nodes represent variables and edges represent direct causal connections. Hard interventions in SCMs are modelled through the do-operator (Pearl, 1995; 2009), which modifies the model’s structure by fixing an endogenous variable to a value $\kappa\in\mathbb{R}$, breaking its original causal dependencies. Formally, applying $do(v_{i}=\kappa)$ to a model $\mathcal{M}$ results in a new model $\mathcal{M}_{v_{i}=\kappa}$ which is identical to $\mathcal{M}$ except that the equation for $v_{i}$ is replaced with a constant value $\kappa$, removing all arrows into $v_{i}$ in the DAG.
@@
Concept-based models Concept-based models (Kim et al., 2018; Chen et al., 2020; Yeh et al., 2020; Kim et al., 2023; Barbiero et al., 2023; Oikarinen et al., 2023) are interpretable architectures that explain predictions using high-level information units (i.e., “concepts”). Most follow the Concept Bottleneck Model (CBM) (Koh et al., 2020) approach. Given a sample’s raw features $\mathbf{x}\in X\subseteq\mathbb{R}^{d}$ (e.g., an image’s pixels), a set of $r$ concepts $c_{i}\in C\subseteq\{0,1\}^{r}$ (e.g., “red”, “round”), and a set of $l$ class labels $y_{j}\in Y\subseteq\{0,1\}^{l}$ (e.g., labels “apple” or “tomato”), a CBM estimates the conditional distribution $\prod_{j}p(y_{j}\mid\text{pa}(y_{j})=c_{1},\dots,c_{r};\theta_{f})\prod_{i}p(c_{i}\mid\mathbf{x};\theta_{g})$ where $p(c_{i}\mid\mathbf{x};\theta_{g})$ is a set of independent Bernoulli distributions. This formulation relies on two key CBMs’ assumptions: i) concepts directly cause class predictions, and ii) concepts are conditionally independent of one another. Usually, concept-based models represent a concept $c_{i}$ using its predicted probability $\hat{c}_{i}\in[0,1]$. However, this may degrade task accuracy when concepts are incomplete (Mahinpei et al., 2021). To address this, Concept Embedding Models (CEMs) (Espinosa Zarlenga et al., 2022) use high-dimensional embeddings $\mathbf{\hat{c}}_{i}\in\mathbb{R}^{z}$ to represent concepts alongside their truth degrees $\hat{c}_{i}\in[0,1]$.
@@ 3 Causal Concept Graph Models
Problem: This work addresses causal opacity, the challenge of uncovering a DL model’s hidden causal structure. Research question: To overcome this, we aim to design DL models where each prediction can be traced back to a chain of semantically meaningful causes. Challenged assumptions: In pursuing this goal, we challenge two key assumptions of Concept Bottleneck Models (CBMs): i) humans know a priori which interpretable variables (concepts) are direct causes of a class label (and datasets embed this prior knowledge), and ii) all interpretable variables used for explanations (i.e., concepts) are conditionally independent. Solution: We introduce Causal Concept Graph Models (Causal CGMs), a class of models making class predictions through cause-effect chains of interpretable variables. We first introduce Causal CGMs, then show how their decision-making process is causally transparent and conclude by describing their learning process.
@@ 3.1 Blueprint
Dropping the assumptions of CBMs—that concepts are independent and direct causes of class labels—requires a model able to capture all possible dependencies [^3] This naive formulation leads to a cyclic probabilistic model, making computation intractable due to the difficulty of handling cycles. Causal Concept Graph Models address this by ensuring efficient learning and producing simple explanations by design (see Sec. 3.3). between interpretable variables (i.e., any classification label). To model such dependencies, Concept Graph Models duplicate the layer of interpretable variables $\mathcal{V}$ by introducing an additional layer of identical copies $\mathcal{V}^{\prime}$ [^4] A detailed analysis of this transformation, including the “unfolding” of the original cyclic model, is provided in App. B and a neural parametrization in App. D.. Using this additional layer, Causal CGMs make predictions of each $v_{i}\in\mathcal{V}$ using as possible inputs all $v_{j}^{\prime}\in\mathcal{V}^{\prime}$ for all $j\neq i$. This inference structure enables CGMs to trace each prediction in $\mathcal{V}$ back to a set of interpretable variables in $\mathcal{V}^{\prime}$. As a result, the set of copies $\mathcal{V}^{\prime}$ in a Causal CGM play the role of “explaining variables”, akin to a CBM’s concepts, while the set of variables $\mathcal{V}$ play the role of “explained variables”, akin to a CBM’s tasks.
@@ 3.2 Causal transparency in Concept Graph Models
The above theorem shows that CGM’s bipartite graph $\mathcal{G}^{\prime}$ between nodes $\mathcal{V}^{\prime}$ and $\mathcal{V}$ can efficiently represent arbitrary long cause-effect chains among interpretable variables (example in Fig. 2). While during training, CGMs exploit the efficient bipartite representation, at test time CGMs can be unfolded following the graph as described in Theorem 3.2 (Sec. 3.3 describes how to learn sparse and acyclic dependencies to make the test time unfolding more efficient and interpretable; we provide more details in App. H.6). This property makes CGMs causally transparent by design, as they can make class predictions through cause-effect chains of endogenous variables, allowing humans to trace each prediction back to its underlying causes within an interpretable parental subgraph. Notice how—unlike CBMs—Causal CGMs do not require to know in advance cause (concept)-effect (task) relationships, nor require that all causes (concepts) are independent from each other, as cause-effect dependencies among interpretable variables are not given, but learnt during training.
@@ 3.3 Training Causal CGMs
Learning sparse, directed, and acyclic structures In order to construct simple explanations, we can make Causal CGMs’ graphs sparse and acyclic. To this end, we parametrise the bipartite graph $\mathcal{G}^{\prime}$ with an adjacency matrix of learnable weights $M\in\mathbb{R}^{k\times k}$. We initialise the weights $m_{ij}\in M$ based on the conditional entropy between labels $v_{i}$ and $v_{j}$ to account for asymmetric concept dependencies (for other initialisation strategies, see App. F). These weights are then fine-tuned through an end-to-end learning process within the Causal CGM framework following common causal priors, where causal graphs are assumed to be sparse, directed, and acyclic (forming a Directed Acyclic Graph or DAG). We introduce a parameter $\gamma\in\mathbb{R}$ to eliminate less significant dependencies and a loss function, as described by Yang et al. (2020), to enforce the sparsity and acyclicity of the causal graph, ensuring that the adjacency matrix $A$ effectively represents a DAG:
@@
Optimisation problem We can now state the general learning objective for Causal CGMs. Given (1) a set of entities represented by their feature vectors $\mathbf{x}\in\mathcal{X}\subseteq X$ (i.e. the input) and (2) a set of annotations for each exogenous variable $v\in\mathcal{V}\subseteq V$ (i.e. the labels), we wish to find functions $\zeta$, $s$, $f$, together with the adjacency matrix $A$, that maximise the log-likelihood of $v,v^{\prime}$, while observing $x$ (or equivalently $u=\zeta(x)$):
@@
where $\lambda_{1,2}$ are hyperparameters balancing optimisation objectives (see related ablation studies in App. H.3). Notice that we use the $do$-operator replacing $\hat{v}_{j}^{\prime}$ with labels $v_{j}$ to reduce leakage and provide better gradients to the endogenous predictor $f$. This enables Causal CGM to be aware of $do$-operations during training, making it effective in responding to $do$-interventions once deployed.
3.4 High-dimensional endogenous representation
As observed by Mahinpei et al. (2021), using scalar representations for concepts (corresponding to endogenous copies in our context) can significantly degrade predictive performance in realistic settings. To address this issue, and inspired by Espinosa Zarlenga et al. (2022), Causal CGMs employ high-dimensional representations for endogenous copies $\mathbf{v}_{j}^{\prime}$. For each endogenous copy, Causal CGMs learn a mixture of two embeddings with explicit semantics, representing the variable’s state. This design enables the model to construct evidence both in favour of and against a variable’s state and supports simple concept interventions, as it allows switching between the two embedding states during interventions. Specifically, we represent the context of each variable with two embeddings: $[\mathbf{c}_{j}^{+},\mathbf{c}_{j}^{-}]=\mathbf{u}_{j}=\zeta(\mathbf{x}),\quad\mathbf{c}_{j}^{+},\mathbf{c}_{j}^{-}\in\mathbb{R}^{m}$. Each embedding carries specific semantics: $\mathbf{c}_{j}^{+}$ represents the active state of the variable, while $\mathbf{c}_{j}^{-}$ represents the inactive state. Once these semantic embeddings are computed, the final endogenous embedding $\mathbf{v}_{j}^{\prime}$ for $v_{j}$ is constructed as a mixture of $\mathbf{c}_{j}^{+}$ and $\mathbf{c}_{j}^{-}$, weighted by the endogenous state:
$\mathbf{v}_{j}^{\prime}=v_{j}^{\prime}\mathbf{c}_{j}^{+}+(1-v_{j}^{\prime})\mathbf{c}_{j}^{-}$ (5)
[…]
@@ 3.5 Causal reasoning and verification with Causal CGM
Ground-truth interventions Causal CGMs support “ground-truth interventions” (see Figure 3(b)). Ground-truth interventions are one of the core motivations behind CBMs (Koh et al., 2020). Through ground-truth interventions, concept bottleneck models allow experts to improve a CBM’s task performance by rectifying mispredicted concepts at test time, thus significantly improving task performance within a human-in-the-loop setting. In Causal CGMs, however, ground-truth interventions have a potential impact on all endogenous variables descendant of an intervened node, which may include not only nodes corresponding to downstream tasks but also nodes corresponding to a CBM’s intermediate concepts. This enables a single concept ground-truth intervention to potentially improve the prediction of intermediate concepts as well as downstream tasks.
Causal reasoning and verification: do-interventions, counterfactuals, and blocking Causal CGMs can answer interventional and counterfactual queries related to the model’s decision-making process using the do-operator on the unfolded SCM.
@@
Figure 3: (a) A 5-variable causal graph. (b) A ground-truth intervention fixes the error of the prediction $\hat{v_{3}}$ to the ground-truth label $v_{3}$. (c) A do-intervention sets the value of the second variable to a constant i.e., $v_{2}=0$. The intervention impacts $v_{2}$’s effects i.e., $v_{3,5}$, but does not alter $v_{2}$’s causes i.e., $v_{1}$. This operation can override ground-truth interventions. (d) A do-intervention on $v_{3}$ blocks the causal effects of $v_{2}$ on $v_{5}$. As a result, intervening on $v_{2}$ cannot alter $v_{5}$ anymore.
@@ 4.1 Key findings
Causal CGMs match the performance of causally opaque models. (Table 1) Causal CGMs demonstrate robust generalisation across all datasets, yielding a predictive performance close to that of black-box architectures with an equivalent capacity. Causal CGMs using a pre-trained causal graph (Causal CGM+CD) tend to have slightly better label accuracy with respect to Causal CGMs where the causal graph is learned end-to-end (Causal CGM). Causal CGMs’s low variance suggests a consistent robustness with respect to weight initialisations over multiple training runs. Thanks to concept embeddings, concept incompleteness settings do not have a strong impact on Causal CGMs’ performance (results in App. H.4).
@@
Ground-truth interventions on Causal CGMs improve both concept and task accuracy as opposed to CBMs (Figures 5, 5) In Causal CGM, the causal graph induces a natural strategy for ground-truth interventions. Indeed, the causal graph narrows down the set of variables to intervene upon: for any given node, we can just fix mispredicted labels of the node’s ancestors as intervening on other nodes will not have any impact. This property significantly decreases the required number of interventions to achieve a desired outcome (e.g., to increase a downstream task accuracy), as shown in App. H. Another advantage of Causal CGM consists in the hierarchical nature of inference that allows ground-truth interventions to impact all endogenous variables descendant of an intervened node. In particular, ground-truth interventions may affect not only nodes corresponding to downstream tasks (as in CBM and CEM), but also nodes corresponding to a CBM’s intermediate concepts. We experimentally verify this property and its impact by calculating—for nodes that were not intervened upon (including both concepts and tasks)—the change in accuracy before and after ground-truth interventions were applied on their ancestors.
@@
Our results (Figure 5 and App. H.1) show that Causal CGM improves nodes accuracy by $\sim 15$ percentage points after only 7 ground-truth interventions on CelebA. CBM and CEM, instead, achieve a similar performance only after intervening on all concepts as their architecture assumes all concepts to be mutually independent. Focusing solely on concept-level analysis, Causal CGM enhances concept accuracy by up to $\sim 25$ percentage points, compared to CBM and CEM, where a concept intervention does not impact the accuracy of other concepts (Figure 5). Causal CGM’s advantage increases with the number of concepts and connections, as a single intervention can impact a higher number of nodes in the causal graph.
Causal CGMs’ endogenous predictors are causally interpretable (Figures 6(a), 6(b)) In Causal CGM the decision-making process is causally interpretable and can be analysed by visualising the learnt causal graph and structural equations as in a structural causal model, as shown in Figure 6(a) for CelebA and in Figure  6(b) for Cifar10 (more results in App. H.2). The first image shows how Causal CGM exploited known biases in CelebA to infer facial attributes. For instance, Causal CGM predicts the attribute “wearing lipstick” directly from the attribute “attractive” and indirectly from attributes such as “smiling” or “high cheek”, all attributes that are known to be strongly correlated with each other in CelebA (Ramaswamy et al., 2021; Wang & Russakovsky, 2023). Similarly, in the CIFAR-10 dataset, where concepts are automatically extracted, the model learns meaningful relationships between concepts (Figure 6(b)), such as the presence of a “port” implying the presence of a “dock”, or the connection between a “beak” and a “bird”. We can also quantify the strength of the causal dependency between two nodes by computing the probability of necessity and sufficiency (PNS) (Pearl, 2022). In the figure, we represent the PNS w.r.t. the leaf node by colouring each node with a different shade of orange (more PNS of the other datasets in App. I). This shows how, for Causal CGM, the attribute “heavy makeup” has the strongest impact on the leaf node. This high degree of causal transparency allows users to interpret Causal CGM’s inference and can be eventually exploited to identify potential biases, thereby supporting the assessment of the model’s counterfactual fairness (an example of generated counterfactuals in Table  2). As a result, users can intervene directly on the causal structure of the decision-making process and remove biases using do-interventions, as shown in the next paragraph.
@@
Causal CGMs can make two causally-related variables causally independent by blocking all paths between these variables (Table 3) The causal transparency of the proposed approach allows users to modify the model’s decision-making process (e.g., to de-bias the model’s inference) by using do-interventions. In particular, we can make two causally-related variables $i$ and $j$ causally independent by blocking all paths between the cause $i$ and the effect $j$. We experimentally verify this property and its impact by computing the Residual Concept Causal Effect i.e., the ratio between the Concept Causal Effect (CaCE) (Goyal et al., 2020) obtained after and before blocking. The optimal value of this metric is zero, corresponding to perfect causal independence between $i$ and $j$ (i.e., the optimal value for a de-biasing operation). The experimental procedure is reported in App. H.7. The results show that, in Causal CGMs, blocking a variable in the causal graph always yields a perfect Residual Concept Causal Effect of zero across all datasets. In contrast, applying the same procedure in CBMs leads only to a negligible reduction in the average causal effect to $3$ percentage points. CEMs not only fail to reduce the causal effect to zero but, in some cases, even increase the causal influence. These results underscore how Causal CGMs transparency enable users to manipulate the model’s decision-making process to achieve desired outcomes, as opposed to existing CBMs.
Table 3: Residual Concept Causal Effect ($\downarrow$) between causally-related variables having blocked all paths between the two variables with do-interventions on the causal graph. The optimal value is zero corresponding to perfect causal independence. (see App. H.7 for further clarification)Values above $100\%$ mean that the causal effect increased instead of decreasing.
checkmark dsprites celeba cifar10